A particular guitar wire is 30.0 cm long and vibrates at a frequency of 196 Hz when no finger is placed on it. The next higher notes on the scale are 220 Hz. 247 Hz, 262 Hz and 294 Hz. How far from the end of the string must the finger be placed to play these notes?

Grade:11

1 Answers

Deepak Patra
9 years ago
Sol. . l base 1 = 30 cm = 0.3 m f base 1 = 196 Hz, f base 2 = 220 Hz We know f ∝ (1/l) (as V is constant for a medium) ⇒ f base 1/f base 2 = l base 1/l base 2 ⇒ l base 2 = 26.7 cm Again f base 3 = 247 Hz ⇒ f base 3/f base 1 = l base 1/l base 3 ⇒ 0.3/ l base 3 ⇒ l base 3 = 0.0224 m = 22.4 cm and l base 3 = 20 cm